Dividend policies and Stock Splits Sophie owns 5,000 shares of a Google-like stock. Which has an Earnings Per Share (EPS) of $100 and a Dividend Per Share (DPS) of $35, and a share price of $1,000. The Google-like stock announces a 10 to 1 stock split. What would be the new EPS, DPS, and share price? How many shares would Sophie have after the split?
Answer #1

Number of shares owned by Sophie= 5000*10 = 50,000

EPS= $100/10 = $10

DPS= $35/10 = $3.5

Share price= $1000/10 = $100